OpenStudy (anonymous):

NEED BIG HELP ON THIS! 10. What is the solution to the rational equation 4x/3 divided by 1/3=5? 11. After x+2/3-x+1/2x have been combined using LCD of 6x, what is the numerator? 12. What is the simplified form of 5/y^2-4/3y 13. What is the simplified form of 8x^4/12y^3 • 20y^8/6x^6 14. What is the simplified form of (x^2-1/9x)/(x^2+2x+1/3x^2) 15. what is the simplified form of x^2-16/x+4 16. what is the soulution to the rational equation (3/2)-(1/x-5)=(-2/2x-10) 17. what is the simplified form of (x/3+y/4)/(x/4-y/3) 18. what is the LCD of (7/3x)+(2/x)+(x/x-1)

OpenStudy (anonymous):

19. what is the simplified form of 4z^2+16z+15/2z^2-z-15 20. what is the siplifed form of y^3+2y-8/y^2+7y+12

OpenStudy (callisto):

10. \[\frac{\frac{4x}{3}}{\frac{1}{3}}=5\]\[\frac{4x}{3}={\frac{5}{3}}\]\[4x=5\]\[x=?\]

OpenStudy (callisto):

11. \[\frac{x+2}{3}-\frac{x+1}{2x}\] \[=\frac{(x+2)(2x)}{3(2x)}-\frac{3(x+1)}{3(2x)}\] \[=\frac{(x+2)(2x)-3x-3}{3(2x)}\] \[=\frac{2x^2+4x-3x-3}{3(2x)}\] \[=\frac{2x^2+x-3}{6x}\] So, what's the answer?

OpenStudy (callisto):

\[\frac{5}{y^2}-\frac{4}{3y}\]\[=\frac{15}{3y^2}-\frac{4y}{3y^2}\]\[=\frac{15-4y}{3y^2}\]

OpenStudy (callisto):

The above one is 12

OpenStudy (callisto):

13 \[\frac{8x^4}{12y^3} •\frac{ 20y^8}{6x^6}\]\[=\frac{2x^4}{3y^3} •\frac{ 10y^8}{3x^6}\]\[=\frac{2(10)y^5}{3(3)x^2} \]\[=\frac{20y^5}{9x^2}\]

OpenStudy (callisto):

I think you need to try the others :|

OpenStudy (anonymous):

what about number 16